OpenStudy (anonymous):

Find the volume of a cone with radius 30 in and slant height 50 in. A. 3769.91 in3 B. 3768 in3 C. 113040 in3 D. 37699.11 in3

OpenStudy (anonymous):



OpenStudy (btaylor):

Volume = 1/3 x π x r^2 x h based on @artix_17 's drawing, h=40. Can you do the rest?

OpenStudy (anonymous):

I don't get it at all...

OpenStudy (anonymous):

volume of cone = 1/3*pi*r^2*h r= 30 h= sqrt (50^2-30^2) =40 1/3*pi*30*30*40 = answer D

OpenStudy (anonymous):

why are these ^ there?

OpenStudy (btaylor):